# Large Language Model Confidence Estimation via Black-Box Access

### v1

@@ Abstract
Estimating uncertainty or confidence in the responses of a model can be significant in evaluating trust not only in the responses, but also in the model as a whole. In this paper, we explore the problem of estimating confidence for responses of large language models (LLMs) with simply black-box or query access to them. We propose a simple and extensible framework where, we engineer novel features and train a (interpretable) model (viz. logistic regression) on these features to estimate the confidence. We empirically demonstrate that our simple framework is effective in estimating confidence of flan-ul2, llama-13b and mistral-7b with it consistently outperforming existing black-box confidence estimation approaches on benchmark datasets such as TriviaQA, SQuAD, CoQA and Natural Questions by even over $10\%$ (on AUROC) in some cases. Additionally, our interpretable approach provides insight into features that are predictive of confidence, leading to the interesting and useful discovery that our confidence models built for one LLM generalize zero-shot across others on a given dataset.
@@ 1 Introduction
Figure 1: Above we see our (extensible) framework to estimate confidence of LLM responses. We propose six prompt perturbations which then can be converted to features based on semantic diversity in the responses and syntactic similarity. The input (tokenized) prompt can optionally be also passed as a feature. The output labels for each (input) prompt are created by checking if the LLM output is correct or not. A (interpretable) logistic regression model is then trained on these features and outputs so that for any new input prompt and LLM response we can estimate the confidence of it being correct based on our model. Moreover, we can also ascertain the features important in estimating these confidences.
@@ 3.1 Prompt Perturbations
Sentence Permutation (SP): If the input has several named entities, we noticed that when the order of the named entities is changed without changing the meaning of the sentence, the output of the model also varied. We first use named entity detector to identify the named entities and then randomly reorder certain number of sentences containing the named entities in the input. Most such reorderings should not affect the output the LLM if it is confident. An example of this is seen in Table 1 third row, where the last sentence in the prompt is now the second sentence.
@@
Stopword Removal (SR): We remove stopwords from the context as specified by the NLTK library. Stopwords are commonly occurring words (viz. "the", "are", "to", etc.) that are assumed to have limited context specific information. Removal of such words should ideally not alter the response of an LLM if the LLM is certain of the answer. An example of this is seen in Table 1 fifth row, where the stopwords are striked out.
@@ 3.2 Featurization
Now based on the above strategies we can construct features to train our confidence model. For each of the first five strategies above we create two types of features: i) based on semantics of the outputs and ii) based on syntactic overlap. For the SRC these are not relevant so we create a different feature as seen below.
@@
Syntactic Similarity: We compute the average syntactic similarity for outputs of each of the first five strategies (run multiple times). The similarity can be measured using standard NLP metrics such as rouge, blue score etc. The higher the syntactic similarity higher the estimated confidence. We use rouge score to quantify the syntactic similarity.
@@
Note that optionally one can also pass the entire prompt as a feature in addition to the above. In the experiments, we saw minimal improvement with such an addition. Semantic set and syntactic similarity were first used by [14] where they applied it only for SD perturbation discussed in the previous section.
@@ 3.3 Label Creation and Confidence Estimation
Table 2: AUROC for the different methods on the four datasets using llama, flan-ul2 and mistral as the LLMs. Higher values indicate better perofmance and the best results are bolded.
@@
Table 3: AUARC for the different methods on the three datasets using llama, flan-ul2 and mistral as the LLMs. Higher values indicate better performance and the best results are bolded.
@@ 4 Experiments
We evaluated the efficacy of our method on Mistral-7B-Instruct-v0.2 [9], llama-2-13b chat version [34], and flan-ul2 models [31]. We elicited responses from these models on four datasets, namely, CoQA [28], SQuAD [27], TriviaQA [12] and Natural Questions (NQ) [16]. CoQA and SQuAD provide the context and expect the model to respond to the question based on the context, while TriviaQA and NQ do not have a context and require the model to tap into its learnt knowledge. For our experiments, we use the validation splits for all the datasets. CoQA has 7983 datapoints, TriviaQA has 9960 datapoints, SQuAD has 10,600 datapoints and NQ has 7830 datapoints.
We used 1000 datapoints to train the Logistic Regression Classifier and the rest of them were used for evaluation. For each of the prompt perturbations specified above, we use five generations for each perturbation. All results are averaged over three runs. We use zero-shot prompting for the datasets with context. For TriviaQA, flan-ul2 and Mistral-7B-Instruct-v0.2 also worked well with zero shot prompting while llama-2-13b chat was performant with a two-shot prompt. For NQ, we used a five shot prompt. The details about the prompts used are provided in the appendix A We used internally hosted models to generate the responses. Thus, we used V100s GPUs for the feature extraction step once the responses were generated. The logistic regression model was trained on an intel core CPU.
We consider methods proposed in recent works [14] and [19] which are state-of-the-art as the baselines. [14] proposed computing the number of semantic sets and lexical similarity metrics from the generated outputs. [19] use eigen value, eccentricity and degree metrics inspired from spectral clustering to estimate the uncertainty of the model. We use the open source code provided by the authors of [19] for comparing with the baselines  [^1] https://github.com/zlin7/UQ-NLG/.
@@
TriviaQA(llama) SD syntactic SD semantic SR syntactic PP syntactic
@@
TriviaQA(flan-ul2) SD syntactic SD semantic PP semantic PP syntactic
@@
TriviaQA(mistral) SD syntactic PP syntactic SP semantic SD semantic
@@
SQuAD(llama) SP syntactic EFA semantic - -
@@
SQuAD(flan-ul2) SP syntactic - - -
@@
SQuAD(mistral) SP syntactic EFA semantic - -
@@
CoQA(llama) SD syntactic EFA semantic SD semantic SR syntactic
@@
CoQA(flan-ul2) SD syntactic EFA semantic SD semantic SP syntactic
@@
CoQA(mistral) SD syntactic SD semantic EFA semantic EFA syntactic
@@
NQ(llama) PP syntactic SD semantic SD syntactic SP syntactic
@@
NQ(flan-ul2) SR semantic SD syntactic SP syntactic PP syntactic
@@
NQ(mistral) PP syntactic SD semantic SD syntactic SP syntactic
@@
CoQA llama 0.71 0.79 (flan-ul2) 0.49 (mistral)
flan-ul2 0.80 0.70 (llama) 0.49 (mistral)
mistral 0.49 0.69 (llama) 0.79 (flan-ul2)
@@
4.1 Confidence Estimation
In Table 2, we see that our method consistently outperforms all the baselines on AUROC. For estimating the confidence of llama’s responses on TriviaQA, our model is better than the best baseline by $11$ percentage points. We are also able to estimate the confidence on the SQuAD dataset using mistral by $14$ percentage points better than the closest competitor. Qualitatively similar results are seen for the SQuAD dataset using flan-ul2 (better by $10$ percentage points) and for the CoQA and NQ datasets using llama (better by $15$ and $10$ percentage points respectively).
We believe these improvements can be attributed to our constructed features and our framework in general. Hence, in the next section we try to ascertain which features for which datasets and LLMs played an important role in predicting the confidences accurately. Note that such an analysis (with high confidence) is possible because our trained model is interpretable. We also tried to pass the tokenized input prompt as additional features (maximum length 256) to our logistic model, however, the improvements were minimal at best and in some cases the performance even dropped possibly because of the model overfitting given that there were now 100s of features. Hence, we do not report these results, although passing the input prompt is still a possibility in general.
@@ 4.2 Confidence Model Interpretability and Transferability
Interpretability: We now study which features in our logistic model were instrumental for accurate confidence estimation. In Table 4, we see the top four features for each dataset-LLM combination. Blanks indicate that there were no features at that rank or lower where their logistic coefficient was greater than $1e^{-4}$. As can be seen the simplest feature SD plays a role in many cases. This indicates that variability of output for the same input prompt is a strong indicator of response correctness. Moreover, other features such as SP and EFA are also crucial in ascertaining confidence as seen in particular for the SQuAD dataset. This points to order bias when looking at contexts and brittleness to redundant information being also strong indicators of response accuracy. PP and SR also play a role in some cases, where they are more crucial for datasets with no contexts such as TriviaQA and NQ. This makes sense as the specific question is more important here in the absence of context and hence the absence of also other features such as SP and EFA. Both the syntactic similarity and semantic set featurizations seem to be important in estimating confidence.
Looking across the datasets and LLMs we see an interesting trend. It seems that for a given dataset different LLMs have similar features that appear to be important. For instance, SP syntactic similarity is the top feature for all three LLMs on SQuAD, while EFA based feature also appears for llama and mistral. For TriviaQA, SD and PP appear for all three models. For CoQA, SD and EFA appear. While for NQ, PP and SD appear as important for all the models. This trend points towards an interesting prospect of applying a confidence estimator of one LLM to other LLMs on a given dataset. As such, we could have a universal confidence estimator just built for one of the LLMs that we could apply across others with reasonable assurance. We explore this exciting possibility in the next part.
@@
This suggests that we could apply our approach to one LLM and then use the same confidence model to evaluate responses of other LLMs without having to build individual models for them. It would be interesting to further stress test this hypothesis in the future with more LLMs and datasets. Nonetheless, even in the current setup – of three LLMs and four datasets – this observation is interesting and useful.
@@ 5 Discussion
In summary, we have provided an extensible framework for black-box confidence estimation of LLM responses by proposing novel features that are indicative of response correctness. By building an interpretable logistic regression model based on these features we were able to obtain state-of-the-art performance in estimating confidence on four benchmark datasets (CoQA, SQuAD, NQ and TriviaQA) and using three powerful open source LLMs (llama-2-13b-chat, Mistral-7B-Instruct-v0.2 and flan-ul2). The interpretability of our confidence model aided in identifying features (viz. SD, SP, EFA,PP) that were instrumental in driving its performance for different LLM-dataset combinations. This led to the interesting realization that many of the features crucial for performance were shared across the confidence models of different LLMs for a dataset. We thus tested if the confidence models generalized across LLMs for a dataset and found that it indeed was the case leading to the interesting possibility of having an universal confidence model trained on just a single LLMs responses, but applied across many others.
Limitations and Broader Impact: Owing to the supervised nature of training the confidence model, one limitation of our approach is that at least some of the model’s generations must be close to the ground truth for us to obtain a reasonable confidence estimator. Another limitation is that the results and insights were obtained based on datasets in English, but these insights might vary when looking at datasets in other languages. More varied tasks and models could be tested upon in the future. We used rouge to test accuracy of generations consistent with previous works, however, rouge, like also other NLP metrics, can be error prone. In terms of broader impact, our approach can be widely applied as it is simple and works with just black-box access to the LLM. Access to logits or internals of the model are not required. However, our estimates although accurate can be imperfect and this should be taken into account when using our approach in high stakes applications involving LLMs. One should also be cognizant of adversaries aware of our features trying to induce misplaced trust in LLMs they create or prefer.
Given the extensibility of our framework, in the future, it would be interesting to add more features as LLMs evolve. One class of such features might be those where the correctness of a response is checked through multiple routes. For instance, given a prompt, one could create questions that are related to the given question and context in the prompt that ideally depend on the response to the original question. Then comparing the model’s responses when the related question is asked just based on the context to when it is asked where the original response is appended to the context could be indicative of confidence. That is if the model is confident, both the responses should be the same. Strategies such as these as well as different ones could be employed going forward.

### v2

@@ Abstract
Estimating uncertainty or confidence in the responses of a model can be significant in evaluating trust not only in the responses, but also in the model as a whole. In this paper, we explore the problem of estimating confidence for responses of large language models (LLMs) with simply black-box or query access to them. We propose a simple and extensible framework where, we engineer novel features and train a (interpretable) model (viz. logistic regression) on these features to estimate the confidence. We empirically demonstrate that our simple framework is effective in estimating confidence of Flan-ul2, Llama-13b and Mistral-7b on four benchmark Q&A tasks as well as of Pegasus-large and BART-large on two benchmark summarization tasks with it surpassing baselines by even over $10\%$ (on AUROC) in some cases. Additionally, our interpretable approach provides insight into features that are predictive of confidence, leading to the interesting and useful discovery that our confidence models built for one LLM generalize zero-shot across others on a given dataset.
@@ 1 Introduction
Figure 1: Above we see our (extensible) framework to estimate confidence of LLM responses. We propose six prompt perturbations which then can be converted to features based on semantic diversity in the responses and lexical similarity. The input (tokenized) prompt can optionally be also passed as a feature. The output labels for each (input) prompt are created by checking if the LLM output is correct or not. A (interpretable) logistic regression model is then trained on these features and outputs so that for any new input prompt and LLM response we can estimate the confidence of it being correct based on our model. Moreover, we can also ascertain the features important in estimating these confidences.
@@ 3.1 Prompt Perturbations
Sentence Permutation (SP): If the input has several named entities, we noticed that when the order of the named entities is changed without changing the meaning of the sentence, the output of the LLM also varied. We first use named entity detector to identify the named entities and then randomly reorder certain number of these sentences. An example of this is seen in Table 1 third row, where the last sentence in the prompt is now the second sentence. As such, if the number of sentences with named entities is less than five, we reorder all of them. If it is greater than five, then we randomly select five and reorder them. Most such reorderings do not affect the LLM output if it is confident.
@@
Stopword Removal (SR): We remove stopwords from the context as specified by the NLTK library. Stopwords are commonly occurring words (viz. "the", "are", "to", etc.) that are assumed to have limited context specific information. Removal of such words should ideally not alter the response of an LLM if the LLM is certain of the answer. An example of this is seen in Table 1 fifth row, where the stopwords are striked out. We ensured that the negative words were not removed as they would change the meaning of the sentence.
@@
As seen in Table 7, the four perturbations above (PP, SP, EFA and SR) that alter the original prompt still maintain the semantics as intended in almost all cases.
@@ 3.2 Featurization
Now based on the above strategies we can construct features to train our confidence model. For each of the first five strategies above we create two types of features: i) based on semantics of the outputs and ii) based on lexical overlap. For the SRC these are not relevant so we create a different feature as seen below.
@@
Lexical Similarity: We compute the average lexical similarity for outputs of each of the first five strategies (run multiple times). The similarity can be measured using standard NLP metrics such as rouge, blue score etc. The higher the lexical similarity higher the estimated confidence. We use rouge score to quantify the lexical similarity.
@@
Note that optionally one can also pass the entire prompt as a feature in addition to the above. In the experiments, we saw minimal improvement with such an addition. Semantic set and lexical similarity were first used by [16] where they applied it only for SD perturbation discussed in the previous section.
@@ 3.3 Label Creation and Confidence Estimation
Table 2: AUROCs on four Q&A and two summarization datasets (CNN, XSUM) using a total of five LLMs (Llama, Flan-ul2, Mistral, Pegasus, BART). Higher values are better. Best results bolded.
@@
CNN (Pegasus) 0.51 0.67 0.73 0.72 0.72 0.77
CNN (BART) 0.51 0.60 0.52 0.48 0.54 0.57
XSUM (Pegasus) 0.51 0.58 0.69 0.70 0.71 0.73
XSUM (BART) 0.51 0.59 0.53 0.51 0.52 0.57
Table 3: AUARCs on four Q&A and two summarization datasets (CNN, XSUM) using a total of five LLMs (Llama, Flan-ul2, Mistral, Pegasus, BART). Higher values are better. Best results bolded.
@@
CNN (Pegasus) 0.45 0.51 0.53 0.43 0.52 0.74
CNN (BART) 0.21 0.22 0.21 0.21 0.21 0.34
XSUM (Pegasus) 0.16 0.17 0.19 0.17 0.17 0.27
XSUM (BART) 0.21 0.22 0.20 0.21 0.22 0.35
@@ 4 Experiments
We demonstrate the efficacy of our method on question answering and summarization tasks. For summarization, we used BART-large [20] and Pegasus-large [46] and for question answering, we used Mistral-7B-Instruct-v0.2 [11], Llama-2-13b chat version [39], and Flan-ul2 models [36]. For question answering we elicited responses from these models on four datasets, namely, CoQA [32], SQuAD [31], TriviaQA [14] and Natural Questions (NQ) [18]. CoQA and SQuAD provide the context and expect the model to respond to the question based on the context, while TriviaQA and NQ do not have a context and require the model to tap into its learnt knowledge. For our experiments, we use the validation splits for all the datasets as done previously [22]. CoQA has 7983 datapoints, TriviaQA has 9960 datapoints, SQuAD has 10,600 datapoints and NQ has 7830 datapoints. For summarization, we used CNN  Daily Mail [33] and [8] and XSUM [27] datasets. We use a subset of the validation splits of both the datasets comprising of 4000 datapoints. For detecting entailment, we use deberta-large-nli model which is specialized for NLI tasks [7].
@@
TriviaQA(Llama) SD lexical SD semantic SR lexical PP lexical
@@
TriviaQA(Flan-ul2) SD lexical SD semantic PP semantic PP lexical
@@
TriviaQA(Mistral) SD lexical PP lexical SP semantic SD semantic
@@
SQuAD(Llama) SP lexical EFA semantic - -
@@
SQuAD(Flan-ul2) SP lexical - - -
@@
SQuAD(Mistral) SP lexical EFA semantic - -
@@
CoQA(Llama) SD lexical EFA semantic SD semantic SR lexical
@@
CoQA(Flan-ul2) SD lexical EFA semantic SD semantic SP lexical
@@
CoQA(Mistral) SD lexical SD semantic EFA semantic EFA lexical
@@
NQ(Llama) PP lexical SD semantic SD lexical SP lexical
@@
NQ(Flan-ul2) SR semantic SD lexical SP lexical PP lexical
@@
NQ(Mistral) PP lexical SD semantic SD lexical SP lexical
@@
CNN(Pegasus) SD lexical EFA lexical SR lexical SP lexical
similarity similarity similarity similarity
CNN(BART) SR lexical SP lexical EFA lexical SP semantic
similarity similarity similarity set
XSUM(Pegasus) SD lexical EFA semantic PP lexical SD semantic
similarity set similarity set
XSUM(BART) SR lexical SP lexical EFA lexical SP semantic
similarity similarity similarity set
We follow previous works [22], which used $1000$ datapoints for hyperparameter tuning, to train our Logistic Regression Classifier and the rest of them were used for evaluation. As such, in Table 8 in the appendix, we show that our method is quite performant even with fewer training datapoints. For each of the prompt perturbations specified above, we use five generations for each perturbation for more robust evaluation. All results are averaged over three runs. We use zero-shot prompting for the datasets with context. For TriviaQA, Flan-ul2 and Mistral-7B-Instruct-v0.2 also worked well with zero shot prompting while Llama-2-13b chat was performant with a two-shot prompt. For NQ, we used a five shot prompt. The details about the prompts used are provided in the Appendix A. We used internally hosted models to generate the responses. Thus, we used V100s GPUs for the feature extraction step once the responses were generated. The logistic regression model was trained on an intel core CPU.
We consider methods proposed in recent works [16] and [22] which are state-of-the-art as the baselines. [16] proposed computing the number of semantic sets and lexical similarity metrics from the generated outputs. [22] use eigen value, eccentricity and degree metrics inspired from spectral clustering to estimate the uncertainty of the model. We use the open source code provided by the authors of [22] for comparing with the baselines  [^1] https://github.com/zlin7/UQ-NLG/ The results are different in some cases from those reported in their paper possibly because of different random splits and different LLMs used, since we did run the provided code..
4.1 Confidence Estimation
In Table 2, we see that our method quite consistently outperforms all baselines on AUROC. For estimating the confidence of Llama’s responses on TriviaQA, our model is better than the best baseline by $11$ percentage points. We are also able to estimate the confidence on the SQuAD dataset using Mistral by $14$ percentage points better than the closest competitor. Qualitatively similar results are seen for the SQuAD dataset using Flan-ul2 (better by $10$ percentage points) and for the CoQA and NQ datasets using Llama (better by $15$ and $10$ percentage points respectively). Our results on the summarization datasets using LLMs that excel at summarization (viz. Pegasus and BART) we see again that we are either better or at least competitive.
We believe these improvements can be attributed to our constructed features and our framework in general. Hence, in the next section we try to ascertain which features for which datasets and LLMs played an important role in predicting the confidences accurately. Note that such an analysis with high confidence is possible because our trained model is interpretable. We also tried to pass the tokenized input prompt as additional features (maximum length 256) to our logistic model, however, the improvements were minimal at best and in some cases the performance even dropped possibly because of the model overfitting given that there were now 100s of features. Hence, we do not report these results, although passing the input prompt is still a possibility in general.
@@ 4.1 Confidence Estimation
CNN Pegasus 0.77 0.57 (BART) -
BART 0.57 0.77 (Pegasus) -
XSUM Pegasus 0.73 0.58 (BART) -
BART 0.57 0.71 (Pegasus) -
@@
CoQA Llama 0.71 0.79 (Flan-ul2) 0.61 (Mistral)
Flan-ul2 0.80 0.70 (Llama) 0.61 (Mistral)
Mistral 0.61 0.69 (Llama) 0.79 (Flan-ul2)
@@
CNN Pegasus 0.74 0.34 (BART) -
BART 0.34 0.74 (Pegasus) -
XSUM Pegasus 0.27 0.34 (BART) -
BART 0.35 0.25 (Pegasus) -
@@ 4.2 Confidence Model Interpretability and Transferability
Interpretability: We now study which features in our logistic model were instrumental for accurate confidence estimation. In Table 4, we see the top four features for each dataset-LLM combination. Blanks indicate that there were no features at that rank or lower where their logistic coefficient was greater than $1e^{-4}$. As can be seen the simplest feature SD plays a role in many cases. This indicates that variability of output for the same input prompt is a strong indicator of response correctness. Moreover, other features such as SP and EFA are also crucial in ascertaining confidence as seen in particular for the SQuAD dataset as well as the summarization datasets. This points to order bias when looking at contexts and brittleness to redundant information being also strong indicators of response accuracy. PP and SR also play a role in some cases, where they are more crucial for datasets with no contexts such as TriviaQA and NQ. This makes sense as the specific question is more important here in the absence of context and hence the absence of also other features such as SP and EFA. Both the lexical similarity and semantic set featurizations seem to be important in estimating confidence.
Looking across the datasets and LLMs we see an interesting trend. It seems that for a given dataset different LLMs have similar features that appear to be important. For instance, SP lexical similarity is the top feature for all three LLMs on SQuAD, while EFA based feature also appears for Llama and Mistral. For TriviaQA, SD and PP appear for all three models. For CoQA, SD and EFA appear. While for NQ, PP and SD appear as important for all the models. This trend points towards an interesting prospect of applying a confidence estimator of one LLM to other LLMs on a given dataset. As such, we could have a universal confidence estimator just built for one of the LLMs that we could apply across others with reasonable assurance. We explore this exciting possibility in the next part.
@@
This suggests that we could apply our approach to one LLM and then use the same confidence model to evaluate responses of other LLMs without having to build individual models for them. It would be interesting to further stress test this hypothesis in the future with more LLMs and datasets. Nonetheless, even in the current setup – of five LLMs and six datasets – this observation is interesting and useful.
@@ 5 Discussion
In summary, we have provided an extensible framework for black-box confidence estimation of LLM responses by proposing novel features that are indicative of response correctness. By building an interpretable logistic regression model based on these features we were able to obtain state-of-the-art performance in estimating confidence on six benchmark datasets (CoQA, SQuAD, NQ, TriviaQA, CNN Daily and XSUM) and using five powerful open source LLMs (Llama-2-13b-chat, Mistral-7B-Instruct-v0.2, Flan-ul2, Pegasus-large and BART-large). The interpretability of our confidence model aided in identifying features (viz. SD, SP, EFA,PP) that were instrumental in driving its performance for different LLM-dataset combinations. This led to the interesting realization that many of the features crucial for performance were shared across the confidence models of different LLMs for a dataset. We thus tested if the confidence models generalized across LLMs for a dataset and found that it indeed was the case leading to the interesting possibility of having an universal confidence model trained on just a single LLMs responses, but applied across many others.
Owing to the supervised nature of training the confidence model, one limitation of our approach is that at least some of the model’s generations must be close to the ground truth for us to obtain a reasonable confidence estimator. Another limitation is that the results and insights were obtained based on datasets in English, but these insights might vary when looking at datasets in other languages. More varied tasks and models could be tested upon in the future. We used rouge to test accuracy of generations consistent with previous works, however, rouge, like also other NLP metrics, can be error prone. In terms of broader impact, our approach can be widely applied as it is simple and works with just black-box access to the LLM. Access to logits or internals of the model are not required. However, our estimates although accurate can be imperfect and this should be taken into account when using our approach in high stakes applications involving LLMs. One should also be cognizant of adversaries aware of our features trying to induce misplaced trust in LLMs they create or prefer.
Given the extensibility of our framework, in the future, it would be interesting to add more features as LLMs evolve. One class of such features might be those where the correctness of a response is checked through creating questions that are (causally) related to the original question and context, and seeing how the response varies by asking this question by itself as opposed to in conjunction with the original question and response. Such and other strategies may help in generalizing these confidence estimators also across datasets something that has been seen when we have additional access to logits of LLMs.